OpenStudy (anonymous):

Solve the triangle. A = 46°, a = 34, b = 27

OpenStudy (anonymous):

B = 34.8°, C = 99.2°, c ≈ 28

OpenStudy (campbell_st):

well you can find B using the law of sines and then C by angle sum of a triangle and c by law of sines or law of cosines

OpenStudy (campbell_st):

well c seems a little out... if C is the largest angle 99 then c has to be the longest side... B seems fine

hero (hero):

@melacho It's best to post only one question at a time.

hero (hero):

Are you here?

hero (hero):

If you want help, you should respond properly. Open Study doesn't work like Yahoo Answers where you just sit and wait for someone to answer your question. We have an expectation for the user to engage in the process of finding the solution to a problem.

OpenStudy (anonymous):

UMMM I know how to do this... and i know everyone that helps me knows i hate when people give me the answers

OpenStudy (anonymous):

@campbell_st B = 34.8°, C = 99.2°, c ≈ 46.7

hero (hero):

Did you figure out the first one yet?

OpenStudy (anonymous):

i only asked one question dude...

OpenStudy (campbell_st):

the answer for c now makes more sense...

OpenStudy (anonymous):

thank you c:

hero (hero):

I think there are two solutions for this.

hero (hero):

@melacho

OpenStudy (anonymous):

care to explain?

hero (hero):

Yes. Basically, I set up the problem using law of sines as such: \[\frac{\sin(46)}{34}= \frac{x}{27}\] \[x \approx 0.57\]

hero (hero):

Of course x = \(\sin B\) so \(\sin B = 0.57\)

hero (hero):

If you take the inverse sine of both sides you get \(\sin^{-1}(0.57) = 34.7502\) However, you should always perform this operation 180 - 34.7502

hero (hero):

Afterwards you will get 146 Then you should check to make sure \(\sin(146)\) isn't also 0.57

hero (hero):

If it is, then that means you have a two triangle case.

hero (hero):

Actually it's more like \(\sin(145.2497) \approx 0.57\)

hero (hero):

Either way, as you can see, you have a two triangle case, so you have to solve for both.

OpenStudy (campbell_st):

but if you use the ambiguous case that can occur in the law of sines... given 1 angle is A = 46 how can a 2nd angle be 146 so the 2nd triangle case doesn't exist..

hero (hero):

I knew there was something else I was supposed to check regarding that. It's been a while.